Perform the following additions and/or subtractions:
(4y^3+8y^2+6y)+(3y^2+2y+9)=

Answer by lenny460(1073)

Perform the following additions and/or subtractions:
(4y^3+8y^2+6y)+(3y^2+2y+9)=

(4y^3 + 8y^2 + 6y) + (3y^2 + 2y + 9)

Remove the parentheses (brackets) and
combine like terms:

8y^2 + 3y^2 = 11y^2
6y + 2y = 8y

Answer:

4y^3 + 11y^2 + 8y + 9




Lennox Obuong
Algebra Tutor
Nairobi, Kenya